What is the percent yield for NO in a reaction with 4 moles of O2? The actual mass obtained from the reaction is 90 grams.
1
Expert's answer
2020-03-24T14:01:41-0400

Assuming nitrogen in excess,

"N_2+O_2\to 2NO"

One mole of "O_2" produce "2" moles of "NO" .So,"4" moles of oxygen will produce "8" moles of "NO" .

Theoretical yield of "NO=8\times (14+16)=240\ g"

Experimental yield "=90\ g"

Percent yield"=\frac{90}{240}\times 100=37.5" %